OpenStudy (anonymous):

if a|(b-1) and a|(c-1) then a|(bc-1)

OpenStudy (theeric):

What is the '|', if I may ask?

OpenStudy (freckles):

divides

OpenStudy (freckles):

like 5|10 because 5(2)=10

OpenStudy (freckles):

5|6 is not a true statement because there is no such integer a so that we have 5a=6

OpenStudy (theeric):

Oh, yeah! :) So you have to work out the proof, I suppose?

OpenStudy (freckles):

yep So you get to play with the following: ak=b-1 where k is an integer ai=c-1 where i is an integer You use the above to show: ap=bc-1 where p is an integer.

OpenStudy (freckles):

you get to play with what you are given to show the then part have you tried doing any playing with those expressions?

OpenStudy (anonymous):

I used the first two equations that you used and i substituted them into the formula a|(ab-1) to obtain a|(a^2ki + ak+ai +1 - 1 but I don't think that is the correct approach

OpenStudy (freckles):

it looks like a great approach

OpenStudy (zarkon):

look at (b-1)(c-1)

OpenStudy (freckles):

and that one a is suppose to be a c but yeah

OpenStudy (anonymous):

so if i end up with a|a(aki + k + i) then that is correct?

OpenStudy (freckles):

well on sec this is the way I would say bc-1=(a(aki+k+i)) so this means a|(bc-1) since aki+j+i is an integer

OpenStudy (anonymous):

that makes a lot more sense than what i put

OpenStudy (freckles):

but your approach was good

OpenStudy (zarkon):

(b-1)(c-1)=bc-c-b+1=bc-1-(c-1)-(b-1) so (b-1)(c-1)+(c-1)+(b-1)=bc-1

OpenStudy (zarkon):

a divides (b-1)(c-1)+(c-1)+(b-1) therefore a divides bc-1

OpenStudy (freckles):

cute stuff zarkson

OpenStudy (anonymous):

thank you both for the help. Appreciate it and you both helped it make a lot more sense

OpenStudy (anonymous):

another question. if a = 7(mod 8) and b = 3(mod 8) then a * b = 1(mod 8). I feel like that is a false statement

OpenStudy (freckles):

well what is the remainder when dividing 21 by 8?

OpenStudy (anonymous):

not 1 that is why i thought it was false

OpenStudy (freckles):

i think you are right :)

OpenStudy (anonymous):

i knew i had a little knowledge of this math stuff haha thanks

OpenStudy (freckles):

np